13 - 2 /-6 + 14 / =
A. -3
B. -27
C. 29
D. 53
Please help ... I am using this sign / as the absolute value... I am very confused .....

Answer by MathLover1(20855)

13 - 2|-6 + 14| = 13 - 2|8| =13-16= -3
A. -3